# Concept Guided Co-saliency Objection Detection

### v1

@@ Abstract
The task of co-saliency object detection (Co-SOD) seeks to identify common, salient objects across a collection of images by examining shared visual features. However, traditional Co-SOD methods often encounter limitations when faced with diverse object variations (e.g., different postures) and irrelevant background elements that introduce noise. To address these challenges, we propose ConceptCoSOD, a novel concept-guided approach that leverages text semantic information to enhance Co-SOD performance by guiding the model to focus on consistent object features. Through rethinking Co-SOD as an (image-text)-to-image task instead of an image-to-image task, ConceptCoSOD first captures shared semantic concepts within an image group and then uses them as guidance for precise object segmentation in complex scenarios. Experimental results on three benchmark datasets and six corruptions reveal that ConceptCoSOD significantly improves detection accuracy, especially in challenging settings with considerable background distractions and object variability.
@@ 1 Introduction
Previous research has addressed the Co-SOD problem from multiple perspectives. Early studies [3, 10, 27] relied on manually selected features, such as color and texture, to identify consistency across related image groups. However, these handcrafted shallow features are limited in robustness and practicality. Consequently, learning-based methods have diversified, incorporating deep learning [26, 39], self-paced learning [17, 46], metric learning [14], and graph-based learning [23, 53] to explore semantic relationships within image groups, achieving substantial advancements. However, two primary challenges remain. The first is inter-object variations, where attributes such as posture vary within the same category (e.g., “dogs”). The second is background distractions, where unrelated elements, like occasional “human” in the background, introduce visual noise. Effective Co-SOD methods must address these challenges by mitigating the influence of both inter-object variations and background distractions.
Figure 1: Co-saliency object detection results w/wo the concept guidance.
With the growing popularity and effectiveness of vision-language methods in segmentation tasks [2, 41, 30, 29, 40], it is worth exploring whether this text-driven approach (i.e., leveraging text semantics for vision tasks) could be adapted to the Co-SOD task. In our opinion, the text-driven approach offers two advantages for addressing the main challenges in the Co-SOD task: ❶ Condensed semantic information in the text (e.g., semantic of “dog”) reduces the impact of subtle variations within the same category, such as posture differences, enabling more precise segmentation of target object regions against inter-object variations. ❷ Under varying visual conditions (e.g., lighting, or background clutter), language provides consistent semantic guidance, allowing the model to focus on the target object despite visual noise from unrelated elements like “human” in the background, thereby enhancing segmentation accuracy and robustness against background distractions.
To this end, instead of treating Co-SOD as a traditional image-to-image task, we propose rethinking it as an (image-text)-to-image task by incorporating text semantics as additional information. Specifically, our method comprises two key modules: ❶ For a given image group containing the same target object, we first extract text-based semantic information of the target object (i.e., concept) related to the target object within the images. These concepts are robust, capturing real-world diversity, including variations in perspective, appearance, and object positioning. ❷ We then leverage these learned concepts to achieve fine-grained segmentation of the target object. We introduce this end-to-end concept-guided co-saliency object detection method as ConceptCoSOD. Fig. 1 illustrates an example of the ConceptCoSOD comparing with traditional co-saliency object detection methods. As shown, traditional co-saliency object detection methods lacking concept guidance produce suboptimal results. In contrast, our concept-guided ConceptCoSOD achieves highly accurate detections on public datasets compared to state-of-the-art methods.
@@
  • To the best of our knowledge, we present the first concept-guided co-saliency object detection method. By fully leveraging text semantics extracted from the input image group as supplementary information, instead of treating Co-SOD as a traditional image-to-image task, we propose rethinking it as an (image-text)-to-image task.
  • We design a novel pipeline consisting of two modules: the concept learning module, which extracts concrete text semantic information, and the concept-guided segmentation module, which leverages the text semantic information to perform fine-grained segmentation directly.
  • Extensive experiments were conducted on three benchmark clean datasets and four datasets with different types of corruptions, confirming the effectiveness and robustness of the proposed ConceptCoSOD method.
@@ 2.1 Co-salient Object Detection
Co-Saliency Detection (Co-SOD) focuses on identifying shared, prominent objects by examining the intrinsic relationships among a group of related images [36, 49]. Traditional Co-SOD techniques [28, 34, 21] relied on handcrafted cues or superpixel-based strategies to capture co-salient areas. However, these manually designed, shallow features often suffer from limitations, proving rigid and insufficient for complex scenarios. With advancements in deep learning, more sophisticated, data-driven Co-SOD approaches have emerged [9, 51], utilizing deeper feature representations to achieve robust consensus formation and feature distribution. These methods first aggregate features from all images in a group to establish a consensus representation, which is then distributed back to the individual image features, enhancing co-saliency extraction. For example, Wang et al. [38] used feature aggregation via summation and applied a gradient feedback mechanism to redistribute the consensus. Jin et al. [24] introduced ICNet, utilizing a set of enhanced intra-saliency vectors and a dense correlation module for precise distribution. Zhang et al. [51] proposed GICD which employs a gradient-induced mechanism that pays more attention to discriminative convolutional kernels that help to locate the co-salient regions. Zhang et al. [48] presented an adaptive graph convolutional network (GCAGC) with attention graph clustering for co-saliency detection. Fan et al. [9] constructed a consensus attention map through an affinity module, which they multiplied back to the individual image features in the proposed method GCoNet. Zhang et al. [50] proposed CADC, encoding consensus information using dynamic kernels and convolving them with image features to effectively distribute consensus knowledge across the group, enhancing overall saliency detection accuracy. Yu et al. [42] introduced a framework called DCFM for mining comprehensive and democratically shared co-salient features without relying on additional information, enhancing detection accuracy.
@@ 2.2 Text-to-Image Diffusion Generation Model
Text-to-image (T2I) generation [44], popularized by diffusion models [4, 16, 31] are highly effective at generating diverse and realistic images based on user-specified text prompts, raising various applications [20]. The development of this approach began with Ho et al. introducing a new generation paradigm, denoising diffusion probabilistic models (DDPMs) [16], which demonstrated competitive performance compared to the widely recognized PGGAN [25]. Recently, the latent diffusion model (LDM) [31] was proposed to perform denoising in latent space, followed by decoding into detailed images, thereby significantly enhancing generation speed and reducing computational complexity. The popular text-to-image model Stable Diffusion [35] is built on LDM.
@@ 2.3 Concept Learning
In concept learning (i.e., personalization) tasks, users provide a few image examples of a desired concept, which are then used to generate novel scenes featuring these newly introduced concepts through text prompts. Current personalization methods typically follow one of two approaches: encapsulating a concept via a word embedding at the input of the text encoder [11, 5], or fine-tuning the weights of diffusion-based modules through various techniques [32, 18, 12, 33, 19]. In our method, we leverage the semantic extraction capability of personalization techniques to capture shared semantics across the image group.
@@ 3.1 Problem Formulation
Given a group of images =$\mathcal{I}=\{\mathbf{I}_{i}\in\mathds{R}^{H\times W\times 3}\}_{i=1}^{N}$ containing $N$ images that have common salient objects, the co-saliency object detection (Co-SOD) task is to detect and segment all common salient objects present within the image group. We denote a Co-SOD method as $\textsc{CoSOD}(\cdot)$ and predict $N$ salient maps via
@@ 3.2 Motivation
Through an investigation of existing Co-SOD methods, we identify two primary challenges.
Inter-object variations. The first challenge arises from significant variations in the shape of the common salient object across different images within the same group. Factors such as posture and viewing angle can lead to inaccurate Co-SOD predictions. For example, in the first row of Fig. 2, the “guitar” is shown from an unusual bottom-up perspective, which causes current state-of-the-art (SOTA) methods to struggle to produce accurate detection results.
Background distractions. The second challenge involves background distractions in group images, which can mislead Co-SOD methods into detecting incorrect objects. As shown in the second row of Fig. 2, the target co-saliency object is the “eggplant” in the bottom left corner of the image. However, existing methods mistakenly shift focus to non-target objects due to these distractions.
Figure 2: Problem of existing Co-SOD methods.
To address these two challenges, our high-level idea comes from the effective practice of vision-language methods in segmentation tasks [2, 41, 29, 40], which exploit text semantic as valuable supplementary information to improve the performance of the task. To be specific, ❶ The condensed semantic information in text minimizes the influence of subtle variations within the same category, such as differences in posture, enabling more accurate segmentation of target object regions amidst inter-object variations. ❷ Under diverse visual conditions (e.g., lighting, or background clutter), language offers consistent semantic guidance, directing the model’s focus toward the target object while mitigating visual noise from unrelated elements, thereby improving segmentation accuracy and robustness against background distractions.
Although leveraging text semantics for the Co-SOD task is promising, no textual information is directly available within the task. Hence, we propose to extract effective semantic information directly from the group of images first and then utilize it to guide subsequent detection. By rethinking the Co-SOD task from an image-to-image to an (image-text)-to-image paradigm, we reformulate the method from Eq. (1) to a new one as follows:
@@
With this design, our approach effectively overcomes the two challenges faced by previous Co-SOD methods in handling difficult scenarios. As shown on the right panel of Fig. 2, our method accurately identifies the correct target object and produces reliable Co-SOD results, regardless of uncommon viewing angles or the presence of distracting background elements.
@@ 4 Methodology: ConceptCoSOD
Figure 3: Group-Image concept learning module.
@@ 4.1 Overview
Figure 4: Visualization of the relationship between the concept learned by our method with image group.
4.2 Group-Image Concept Learning
This module aims to learn the text semantic information (i.e., concept) $\textbf{c}^{*}$ from the image group $\mathcal{I}$. Drawing inspiration from the semantic extraction capabilities of personalization techniques [11] which can generate images of a specific object in diverse scenes using only a few photos of that object, we adopt this approach to capture the concept of the target object from the image group. Note that the concept $\textbf{c}^{*}$, learned by the $\text{ConceptLearn}(\cdot)$ module, represents a token in the latent space of the diffusion model, encapsulating the semantic information of the co-existing object within the image group. The architecture of this module is as below and shown in Fig. 3.
Architecture of text-to-image diffusion model. We utilize a classic text-to-image diffusion model [31] to search for the concepts we need to learn. The text-to-image diffusion model consists of three core modules: (1) an image autoencoder, (2) a text encoder, and (3) a conditional diffusion model. The image autoencoder module comprises an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. The encoder maps the input image $\mathbf{X}$ to a low-dimensional latent space $\textbf{z}=\mathcal{E}(\mathbf{X})$, while the decoder converts the latent representation back to the image space $\mathcal{D}(\mathcal{E}(\mathbf{X}))\approx\mathbf{X}$. The text encoder $\Gamma$ first tokenizes the text y and then translates the tokenized results into text representations in the latent space, $\Gamma(\textbf{y})$. The input to the conditional diffusion model $\epsilon_{\theta}$ consists of the time step $t$, the noise latent variable at step $t$, $\textbf{z}_{t}$, and the text representation $\Gamma(\textbf{y})$. This conditional diffusion model can predict the noise added to $\textbf{z}_{t}$ at time step $t$, denoted as $\epsilon_{\theta}(\textbf{z}_{t},t,\Gamma(\textbf{y}))$.
@@ 4.2 Group-Image Concept Learning
$\textbf{c}^{*}=$ $\operatornamewithlimits{arg\,min}\limits_{\textbf{c}}\mathbb{E}_{\textbf{z},\mathbf{y},\epsilon\in\mathcal{N}(0,1),t\sim\mathcal{U}([0,1])}$
$(\|\epsilon_{\theta}(\mathbf{z}_{t},t,\Phi(\Gamma(\mathbf{y}),\textbf{c}))-\epsilon\|_{2}^{2}),$ (5)
where $\mathbf{y}$ is a fixed text (i.e., ‘a photo of $S^{*}$’), and the function $\Phi(\Gamma(\mathbf{y}),\textbf{c})$ represents an injection process that replaces the ‘$S^{*}$’ token in $\Gamma(\mathbf{y})$ with the learnable embedding c. During the optimization process of this formula, the concept $\mathbf{c}$ is iteratively optimized such that the noise predicted by the diffusion model gradually approaches the ground truth noise $\epsilon$. The Eq. (5) forces the optimized concept $\mathbf{c}^{*}$ to represent the position of the co-salient objects in the image group within the text latent space. After obtaining $\mathbf{c}^{*}$, we can embed the ‘$S^{*}$’ token into other texts and generate new images using the text-to-image diffusion model.
To verify the effectiveness of the concepts, we visualize the concept within the image group in Fig. 4. The results demonstrate that the learned concept is closely aligned with the target object (i.e., axe).
@@
4.3 Concept-Guided Segmentation
4.3.1 Preliminary Exploration
To investigate whether the learned concept $\textbf{c}^{*}$ can be applied to the co-saliency object detection task, we leverage the well-known interpretability method DAAM [37] for a preliminary exploration. DAAM is designed to interpret the internal mechanisms of large-scale text-to-image models. For a given text input and a text-to-image model, DAAM analyzes the associations between individual words in the text and specific elements within the generated image, visualizing how input text prompts influence different regions of the image. It is implemented by exploiting the attention block (containing a self-attention layer and a cross-attention layer) in the UNet portion of the conditional text-to-image model. Specifically, DAAM uses the text representation and latent representation (i.e., the image embedding) as input, extracting the output features (i.e., attention maps) of the cross-attention layer to illustrate the relationships between the latent image representation and the text representation.
Given the DAAM method as a function $\text{AttnExtract}(\cdot)$, We exploit the optimized concept learned from the image set as $\textbf{c}^{*}$, and the process of extracting features from the cross-attention layer is represented as:
$\textbf{f}_{attn}=\text{AttnExtract}(\epsilon_{\theta},\mathbf{z}_{t},t,\textbf{c}^{*}),$ (6)
where $\mathbf{z}_{t}$ represents the latent representation corresponding to the image at time step $t$, $\epsilon_{\theta}$ denotes the parameters of the text-to-image model, and $\textbf{f}_{attn}$ refers to the association features (a set of attention maps) that we extract from the selected cross-attention layers. To obtain the final attention map $\mathbf{S}$, we apply average operation on the $\textbf{f}_{attn}$:
$\mathbf{S}=\text{Avg}(\textbf{f}_{attn}).$ (7)
Fig. 4 demonstrates the effectiveness of DAAM in segmenting the co-salient object. The first row shows images with an axe as the target co-salient object. The second row overlays the predicted co-saliency detection results as heatmaps on the original images for a clearer, more intuitive visualization. The third row presents the final co-saliency object detection results obtained using DAAM. It can be observed that DAAM demonstrates the potential for co-saliency detection, but the visualization results lack precision. Specifically, while the contours of the target axe are discernible in the predicted grayscale images, the segmentation is coarse-grained due to two main disadvantages. ❶ Within the axe target region, certain local areas exhibit relatively low association scores. ❷ Outside the axe region, residual attention is present in areas where ideally no attention should be focused.
To this end, we propose to refine the attention map with a segmentation network to ensure better results in terms of accuracy and detail.
4.3.2 Architecture Design
The concept-guided segmentation module contains two submodules. As shown in Fig. 5, we use an attention extraction submodule to extract the coarse attention map first and then take it as the guidance input for the fine-grained segmentation submodule to output the final segmentation result. Both the submodules are classical U-Net architecture used in text-to-image models whose process takes multi-modality input (refer to architecture [31]).
Specifically, given an image $\mathbf{I}$ and its corresponding concept representation $\textbf{c}^{*}$, the concept-guided segmentation module first transforms the image into a latent representation $\mathbf{z}$ at time step $t$. Using this latent $\mathbf{z}_{t}$ and concept representation $\textbf{c}^{*}$ as input, the attention extraction submodule generates an attention map following Eq. (6). Notably, the fine-grained segmentation submodule differs slightly from the traditional U-Net in text-to-image model by incorporating an additional cross-attention mechanism to process the attention map $\textbf{f}_{attn}$ as supplementary guidance alongside the prior guidance (i.e., concept representation $\textbf{c}^{*}$). With latent $\mathbf{z}_{t}$, concept representation $\textbf{c}^{*}$, and attention map $\textbf{f}_{attn}$ as inputs, the fine-grained segmentation submodule (i.e., function $\text{FineSeg}(\cdot)$) leverages these components to produce the final segmentation output,
@@ 4.3.2 Architecture Design
where $\epsilon_{\phi}$ denotes the parameters in the module. Finally, we apply a predefined threshold value $\lambda$ to $\mathbf{S}^{\prime}$ for final processing. In subsequent experiments, we discuss the impact of choosing different threshold values $\lambda$ on the co-saliency object detection results. For each pixel $s$ in $\mathbf{S}^{\prime}$, the value satisfies the following condition:

### v2

# Concept Guided Co-salient Object Detection

@@ Abstract
Co-salient object detection (Co-SOD) aims to identify common salient objects across a group of related images. While recent methods have made notable progress, they typically rely on low-level visual patterns and lack semantic priors, limiting their detection performance. We propose ConceptCoSOD, a concept-guided framework that introduces high-level semantic knowledge to enhance co-saliency detection. By extracting shared text-based concepts from the input image group, ConceptCoSOD provides semantic guidance that anchors the detection process. To further improve concept quality, we analyze the effect of diffusion timesteps and design a resampling strategy that selects more informative steps for learning robust concepts. This semantic prior, combined with the resampling-enhanced representation, enables accurate and consistent segmentation even in challenging visual conditions. Extensive experiments on three benchmark datasets and five corrupted settings demonstrate that ConceptCoSOD significantly outperforms existing methods in both accuracy and generalization.
@@ 1 Introduction
Previous Co-SOD research explored various approaches. Early methods [3, 10, 27] relied on handcrafted features like color and texture but lacked robustness and scalability. More recent learning-based techniques, such as deep learning [25, 37], self-paced learning [18, 45], metric learning [15], and graph reasoning [22, 51], have significantly improved semantic modeling and overall performance. However, despite these advances, most existing methods overlook the crucial role of semantic priors, revealing a fundamental flaw in their design. They typically approach object discovery as a low-level pattern matching task, relying on visual cues such as color, texture, or spatial layout, without incorporating high-level knowledge about object categories or concepts. Without such priors, the model lacks a conceptual anchor to distinguish essential features from incidental ones. As a result, it struggles with inter-object variations because appearance differences, such as changes in pose or scale, can obscure the underlying semantic identity of the object. Similarly, the absence of semantic grounding makes the model vulnerable to background distractions, where non-target elements that share similar visual patterns with the object can easily mislead detection. In both cases, the model fails not because of insufficient data or inadequate architecture, but because it lacks a high-level understanding of what to look for.
To address these issues, we propose incorporating semantic priors into the Co-SOD pipeline. Specifically, we introduce a concept-guided framework that leverages the rich object-level knowledge encoded in vision-language models such as diffusion-based text-to-image models. Through a personalization process, we extract a concept representation of the target object that captures high-level semantics and structural cues. This representation serves as a semantic anchor that guides the model in identifying the true object of interest, thereby improving robustness to both inter-object variations and background distractions. Our approach consists of two key modules: ❶ Given a group of related images, we first extract text-based semantic information associated with the target object, forming a personalized concept that reflects real-world diversity in pose, appearance, and perspective. ❷ This concept is then used to guide fine-grained segmentation of the target object across the image group. To further improve task performance, we analyze the effect of the concept-learning process on Co-SOD outcomes and identify that the choice of timestep in the diffusion model significantly impacts final detection quality. Based on this observation, we design a novel timestep resampling strategy for more effective object localization. This leads to a clear improvement in overall Co-SOD accuracy and robustness. We term the overall end-to-end framework ConceptCoSOD.
@@
  • We propose the first concept-guided Co-SOD framework that explicitly integrates semantic priors from diffusion models. Unlike previous methods based on low-level pattern matching, our approach uses high-level concepts from the image group to guide co-saliency detection, offering a more principled and semantically informed solution.
  • We propose a two-stage pipeline: a concept learning module to distill object-level semantics and a concept-guided segmentation module for precise object localization.
  • We further introduce a timestep resampling strategy, motivated by an analysis of diffusion timesteps, which enhances concept quality and further boosts Co-SOD performance.
  • Extensive experiments on seven datasets and five baselines demonstrate the effectiveness and robustness of the proposed ConceptCoSOD method.
@@ 2.1 Co-salient Object Detection
Co-Saliency Detection (Co-SOD) identifies shared salient objects in related images by exploiting their intrinsic relationships [35, 46]. Early methods [26, 33, 20] used handcrafted features, which lacked robustness in complex scenes. Recent deep learning approaches [9, 49] improve performance by forming a group-wise consensus representation and redistributing it to refine image saliency.
For instance, Wang et al. [36] used summation-based feature aggregation with gradient feedback. Jin et al. [23] proposed ICNet, incorporating enhanced intra-saliency vectors and dense correlations. Zhang et al. [49] introduced GICD, emphasizing discriminative kernels via gradient induction. Zhang et al. [47] developed GCAGC, an adaptive graph convolutional network with attention-based clustering. Fan et al. [9] built GCoNet using an affinity-based consensus attention map. CADC [48] used dynamic kernels for effective consensus distribution. Lastly, Yu et al. [40] presented DCFM, a democratic feature mining framework enhancing detection without extra inputs.
@@ 2.2 Text-to-Image Diffusion Generation Model
Text-to-image (T2I) generation [42], popularized by diffusion models [4, 17, 30], excels at producing diverse, realistic images from text prompts. This began with Ho et al.’s denoising diffusion probabilistic models (DDPMs) [17], which rivaled PGGAN [24]. Later, the latent diffusion model (LDM) [30] was proposed to perform denoising in latent space, followed by decoding into detailed images, thereby significantly enhancing generation speed and reducing computational complexity. Stable Diffusion [34], a widely used T2I model, is based on LDM.
@@ 2.3 Concept Learning
In concept learning (personalization), users provide a few example images to generate new scenes featuring the desired concept via text prompts. Current personalization methods typically follow one of two approaches: encapsulating a concept via a word embedding at the input of the text encoder [11, 5], or fine-tuning the weights of diffusion-based modules through various techniques [31, 19, 13, 32]. Our approach utilizes these personalization techniques to extract shared semantics across image groups.
@@ 3.1 Problem Formulation
Given a group of images =$\mathcal{I}=\{\mathbf{I}_{i}\in\mathds{R}^{H\times W\times 3}\}_{i=1}^{N}$ containing $N$ images that have common salient objects, the co-saliency object detection (Co-SOD) task is to detect and segment all these objects in the image group. We denote a Co-SOD method as $\textsc{CoSOD}(\cdot)$ and predict $N$ salient maps via
@@ 3.2 Motivation
We investigated existing Co-SOD methods and revealed that they struggle to maintain performance when dealing with complex scenes. The reason is that these Co-SOD methods typically treat the discovery of co-salient objects as a low-level pattern matching task, relying heavily on visual cues such as color, texture, or spatial layout. Although these methods are capable of capturing surface-level similarities among objects, they often lack the ability to model deeper semantic relationships essential for the co-salient object understanding.
As illustrated in Fig.1, existing Co-SOD methods exhibit significant inaccuracies in both presented examples. In Fig.1 (a), the group images contain guitars viewed from different angles. The first image presents a bottom-up perspective, where current state-of-the-art (SOTA) methods consistently fail to detect the entire guitar. Instead, they isolate only the guitar strings and erroneously treat them as the complete object. This error stems from the lack of prior knowledge about the object’s structural composition. Without understanding that a guitar is made up of components such as the body, neck, and headstock, the model misinterprets a partial observation as the whole. In Fig. 1 (b), the co-salient object is an apple, yet the first image also contains a lemon. Without any conceptual understanding of what constitutes an apple, the model fails to recognize the lemon as a separate entity and mistakenly treats it as part of the apple. This further reflects the absence of high-level semantic priors that are necessary for distinguishing object boundaries and identities. These examples reveal a fundamental limitation of prior-free Co-SOD methods. Lacking semantic understanding, these models rely solely on low-level appearance consistency across images and fail to reason about object completeness or category membership. Consequently, they misidentify parts as wholes and incorporate unrelated distractors into the co-salient target, especially under viewpoint variation or background clutter.
Figure 1: Limitations of existing prior-free Co-SOD methods.
Therefore, it is necessary to incorporate prior knowledge into the Co-SOD process. Existing vision-language models have been extensively trained and possess rich semantic knowledge. We aim to leverage their capabilities by incorporating this prior knowledge into our Co-SOD framework. We drew inspiration from the effective practice of vision-language methods in segmentation tasks [2, 39, 28, 38], which exploit text semantics as valuable supplementary information.
Although leveraging text semantics for the Co-SOD task is promising, no textual information is directly available within the task. Hence, we propose to extract effective semantic information directly from the group of images first and then utilize it to guide subsequent detection. Thus, we reformulate the method from Eq. (1) to a new one as follows:
@@ 4.1 Overview
4.2 Concept Learning Module
This module learns the semantic concept $\mathbf{c}^{*}$ from the image group $\mathcal{I}$, inspired by personalization methods [11] that extract object-level concepts from a few images. Specifically, $\mathbf{c}^{*}$ is a learnable token in the latent space of a text-to-image diffusion model, representing the shared object across the group. The module architecture is illustrated in Fig. 2.
Figure 2: Group-Image concept learning module.
We adopt a text-to-image diffusion model [30] to guide concept learning, consisting of three components: (1) an image autoencoder with encoder $\mathcal{E}$, and decoder $\mathcal{D}$, where $\mathcal{E}$ maps an image $\mathbf{X}$ to a latent vector $\mathbf{z}=\mathcal{E}(\mathbf{X})$ and $\mathcal{D}$ can reconstruct it with $\mathcal{D}(\mathcal{E}(\mathbf{X}))\approx\mathbf{X}$, (2) a text encoder $\Gamma$ that converts a text prompt $\mathbf{y}$ into a latent embedding $\Gamma(\mathbf{y})$, and (3) a conditional diffusion model $\epsilon_{\theta}$ that can predict the noise added to $\textbf{z}_{t}$ at time step $t$, denoted as $\epsilon_{\theta}(\textbf{z}_{t},t,\Gamma(\textbf{y}))$, with time step $t$, the noise latent variable at step $t$, $\textbf{z}_{t}$, and the text representation $\Gamma(\textbf{y})$ as input.
@@ 4.2 Concept Learning Module
$\textbf{c}^{*}=$ $\operatornamewithlimits{arg\,min}\limits_{\textbf{c}}\mathbb{E}_{\textbf{z},\mathbf{y},\epsilon\in\mathcal{N}(0,1),t\sim\mathcal{U}([0,1])}(\|\epsilon_{\theta}(\mathbf{z}_{t},t,\Phi(\Gamma(\mathbf{y}),\textbf{c}))-\epsilon\|_{2}^{2}),$ (5)
where $\mathbf{y}$ is a fixed text prompt (e.g., ‘a photo of $S^{*}$’), and $\Phi(\Gamma(\mathbf{y}),\mathbf{c})$ denotes the injection process that replaces the ‘$S^{*}$’ token in $\Gamma(\mathbf{y})$ with the learnable embedding $\mathbf{c}$. Eq. (5) encourages the optimized concept $\mathbf{c}^{*}$ to encode the co-salient object positions in the text latent space. Once $\mathbf{c}^{*}$ is obtained, it can be inserted into other prompts to guide new image generation via a text-to-image diffusion model.
4.3 Concept-Guided Segmentation
The concept-guided segmentation module contains two submodules. As shown in Fig. 3, we use an attention extraction submodule to extract the coarse attention map first and then take it as the guidance input for the fine-grained segmentation submodule to output the final segmentation result. Both the submodules are classical U-Net architectures used in text-to-image models whose process takes multi-modality input (refer to architecture [30]).
@@ 4.3 Concept-Guided Segmentation
Specifically, given an image $\mathbf{I}$ and its corresponding concept representation $\textbf{c}^{*}$, the concept-guided segmentation module first transforms the image into a latent representation $\mathbf{z}$ at time step $t$. Using this latent $\mathbf{z}_{t}$ and concept representation $\textbf{c}^{*}$ as input, the attention extraction submodule generates an attention map following $\textbf{f}_{attn}=\text{AttnExtract}(\epsilon_{\theta},\mathbf{z}_{t},t,\textbf{c}^{*})$. Notably, the fine-grained segmentation submodule differs slightly from the traditional U-Net in text-to-image model by incorporating an additional cross-attention mechanism to process the attention map $\textbf{f}_{attn}$ as supplementary guidance alongside the prior guidance (i.e., concept representation $\textbf{c}^{*}$). With latent $\mathbf{z}_{t}$, concept representation $\textbf{c}^{*}$, and attention map $\textbf{f}_{attn}$ as inputs, the fine-grained segmentation submodule (i.e., function $\text{FineSeg}(\cdot)$) leverages these components to produce the final segmentation output,
@@
where $\epsilon_{\phi}$ denotes the parameters in the module. Finally, we apply a predefined threshold value $\lambda$ to $\mathbf{S}^{\prime}$ for final processing. In subsequent experiments, we discuss the impact of choosing different threshold values $\lambda$ on the co-saliency object detection results. For each pixel $s$ in $\mathbf{S}^{\prime}$, the value satisfies:
@@
4.4 Timestep Resampling Strategy
It is evident that the quality of the concept learned from input images plays a critical role in determining segmentation performance in the CoSOD task. This raises an important question: is the concept representation learned by existing personalization methods sufficiently effective for CoSOD? To investigate this, we compare the segmentation performance achieved using the personalized concept representation $\textbf{c}^{*}$ with that obtained using native token representations (e.g., the textual tokens “cat”, “dog”) directly from the token dictionary of a text-to-image model. As shown in Figure 4 (a), native token-based concepts achieve significantly better segmentation performance than the personalized concept, suggesting that the concept learned through personalization still leaves considerable room for improvement in the context of the CoSOD task.
To improve concept learning from input images in the CoSOD task, we conduct an empirical study on a critical factor in diffusion models: the sampling timestep. As shown in Figure 4 (b), our results demonstrate that the choice of timestep significantly affects the final segmentation performance. Motivated by this observation, we design a timestep resampling strategy to enhance concept learning. Specifically, based on the performance trends in Figure 4 (b), we introduce a resampling method controlled by a ratio parameter $\alpha$. We divide the full diffusion range $T$ into three intervals: the head $T_{1}$, the middle $T_{2}$, and the tail $T_{3}$. The resampling distribution $p_{\alpha}(t)$ is defined as:
$p_{\alpha}(t)=\begin{cases}\frac{1}{2\cdot|T_{1}|\cdot(1+\alpha)},&t\in T_{1}\\ \frac{\alpha}{|T_{2}|\cdot(1+\alpha)},&t\in T_{2}\\ \frac{1}{2\cdot|T_{3}|\cdot(1+\alpha)},&t\in T_{3}\\ \end{cases}$
This formulation ensures that a higher value of $\alpha$ increases the sampling probability in the informative middle interval $T_{2}$, while proportionally decreasing it in the less effective regions $T_{1}$ and $T_{3}$. The distribution is normalized so that the total sampling mass remains one, thereby preserving the overall sampling budget.
Figure 4: Exploring the potential of ConceptCoSOD and the effectiveness of different timestep ranges.
(a) Comparison between the base version of ConceptCoSOD (i.e., without resampling strategy) and the theoretical upper bound (i.e., with accurate tokens). (b) Performance comparison across ten different timestep ranges.